MtKent
Home Obs Webcams
Weather Temp's Motors
Tenerife
Home Status OR Ins.
Weather Temp's Sat Im
Motors Obs Webcams
Find -tasks- here

SONG technical logbook page

Choose site
Logbook for MtKent selected
Project
Logbook entries for all projects are shown
Add to tasks?
Should someone look more into this or is it fixed?
Date of event
Description
Who are you?
Hardware group
Action performed
Attachment
[Maximum size = 20MB]
Notify
Send out notification e-mail? Messages will always be send to slack -> SONG-operations.

Search the log:
DateDescriptionPerson(s)Hardware GroupAction performedAttachmentEditSiteTask(?)
2025-11-27 The flux level of SONG T1 is very low. Mads Telescope 1 It looks like it must be somewhere inside or before the nasmyth guiding unit since not much light hits the guide camera either. Maybe some bug are infesting the guide unit? or something else in the telescope is reducing the amount of light significantly. None Edit MtKent 278
2025-11-12 The boltwood weather data periodically (every few to five minutes) reports values that incorrectly suggest rain. Since you are going out to Mt. Kent, could you please check up on that? Thanks Karsten None Edit MtKent 276
2025-11-12 The door to the electronics cabinet in the SONG dome 2 seems to be open Mads Dome 2 It needs to be closed to prevent hitting the telescope or cables... None Edit MtKent 275
2025-10-27 Cannot connect to telescope in PWI4. Get Error: "2025-10-28 02:15:52.195: ERROR System.Exception: Unable to connect to Azimuth/RA axis: The operation has timed out. at PWI4.MotionControllers.ElmoMotionController.Connect()". This seems similar to earlier error when the power to the motor was somehow disconnected. Please investigate. Karsten None Edit MtKent 273
2025-10-10 Low flux level during some time Mads, Frank Spectrograph Check all motor combinations and figured out that the Iodine cell position has changed....? The new position is now added to the code and all should run correctly. None Edit MtKent
2025-09-19 Kept closed because of controlled forrest fires (fuel reduction burns) around the observatory Mads All Stopped the monitor and closed down Show Edit MtKent
2025-09-16 No power in the FUT dome since 03.50 . Likely breaker switch, since SONG teleskopes also lost power, but that is back Karsten None Edit MtKent 272
2025-09-16 Full power failure at Mt Kent Mads All Had to restart everything since all had been powered off. We are now observing again and all seems fine. Temp. box will slowly stabilize in a day or two. None Edit MtKent
2025-08-28 The SONG telescope 2 reports -Azm Error 29956 Motor pulse usage too high- and will disable motors after short movement. Mads Telescope 2 I have tried to power cycle the telescope control box and restart PWI 2 and the PC. It seems it is and actual issue with the telescope azimuth movement. Not sure if it needs greasing or something else to ease the movement? Show Edit MtKent 271
2025-08-22 Duncan and Frank adjusted the spectrograph Frank Spectrograph No ThAr light: The controller did not deliver a sufficiently high voltage - We switched to the old, non-adjustable power supply. This worked! We adjusted the iodine mirror such that the fiber outputs fall exactly on the ThAr fiber output. All modes (flat, flati2, fp and ThAr) seem to work right now and give light on the detector. None Edit MtKent
2025-08-19 No connection to IRF90 (rotator and focuser). Are power and connections to the unit OK? The values reported for the focus position have been acting strangely for days prior to this. Karsten None Edit MtKent 270
2025-08-06 MADS: For AFTER the vacation ! :-) Ghost in the machine ???? Yesterday I stopped both T1 and T2 monitors and or-scheduler --- yet, I discovered in the afternoon that the telescopes had opened and tried to observe, however acquisition always seemed to fail. In addition at the end of the night, when the sun was 3.5deg below the horizon both domes were still open!! Frank Telescope I Closed the domes manually from HW. One observation + many ThAr was obtained in the night. NO flux on the ThArs. None Edit MtKent 269
2025-08-05 MADS: I put urgent, but guess you are on vacation, so this is for after :-) :The weather is great and not errors reported anywhere - but no observations. Frank Telescope It seems that I cannot home/move T1, T2 using the PWI2 interface. Tried to restart PWI on T1 ... seemingly no effect. I have disabled T1, T2 monitors and the OR_scheduler and closed the domes. None Edit MtKent 268
2025-07-28 No calibration light through fibres for quite some time Mads calibration shutters Checked if light from ThAr and Flat came to guide unit in nasmyth which it did not - most likely the shutters in the guide unit are not activated in remote mode. @Duncan: could you please check and activate them next time at MtKent? None Edit MtKent 266
2025-07-24 Cannot connect to telescope and telescope in strange position. When trying to connect through PWI4 one gets an error: Axis0 communication channel not found. Is the device plugged in and turned on? See attachment. Duncan, feel free to call me at 91959459 at any time if you need help troubleshooting or to test something. Karsten Show Edit MtKent 263
2025-07-07 Disabled usage of T2 since this is causing issues on a daily basis and most people are on holidays.... Mads Telescope 2 Closed down and turned off monitor. None Edit MtKent
2025-06-30 T2 did not home and errors where shown in PWI2 - Tracking error Mads Telescope 2 Re-enabled the motor and then moved the telescope a bit. This raised another error (motor pulse too high). I then power cycled the PW control box. Re-started PWI2, connected and enabled and then the homing went well. Show Edit MtKent
2025-06-24 From the daily light on slit checks it was noticed that light with iodine was not centered on slit Mads, Duncan, Frank, Jonatan Spectrograph The spectrograph box was opened and the Iodine cell was rotated until spots were aligned on slit. Afterwards it was noticed that this was most likely not needed and the confusing was originating from misconfiguration for the daily check images. None Edit MtKent
2025-06-24 Calibration light from fibre in SONG T1 was not centered on science fibre. Mads, Duncan Calibration unit The fibre outout from the calibration fibre was adjusted using the alignment screws. The alignment is not perfect but much better now. None Edit MtKent
2025-06-23 The calibration fibre for SONG telescope 1 is misaligned with the science fibre. See attachment Mads Calibration unit Light from calibration unit was checked on the guide unit in the nasmtyh of T1. Afterwards light was send back through the science fibre to determine the location on the guider. These two do not align and the calibration fibre should be adjusted on site to hit the science fibre better. Show Edit MtKent 260
2025-06-12 The shutter on T1 does not seem to work - probably not in remote mode. Please go to T1 and press the button to activate remote mode for the shutter. Frank Telescope Please also check that the T2 shutter is in remote mode. None Edit MtKent 257
2025-06-10 Re-occurring problem with T2 - Azm error - Motor pulse usage too high Mads, Frank Telescope 2 Powered off the PW box in dome 2, waited some seconds and powered on again. Restarted PWI-2 and moved the telescope a bit with the arrows before homing. Show Edit MtKent
2025-06-06 Telescope reported not ready in web control and on slack and email there was a message that PWI4 was not running Karsten Telescope On remote desktop PWI4 was running, but PWI3 and ASCOM server were not. I started PWI3 and ASCOM server and the checked that the web controls were running again None Edit MtKent
2025-06-06 T2 did not find targets Mads Telescope 2 Did a homing sequence of T2 again and it would then find targets. None Edit MtKent
2025-06-05 The observer page for Mt. Kent shows that all is good - but no signal from T2 - guiding is active, but no star in image and no spectrum from T2 Frank Slit-guider - T2 None, this can wait until next working day. None Edit MtKent 255
2024-07-02 Raw images of ThAr spectra show nothing from 2. July Jonatan Rudrasingam Spectrograph Added it to the task as time worthy, since the wavelength solution on Mt. Kent is stable and people are on holiday. But it needs be looked at somepoint during August at the latest None Edit MtKent 261